OpenStudy (pokemon23):

8b-4(b-2)=24 8b-4(b-2)=24 @Mathematics

OpenStudy (pokemon23):

show all work

OpenStudy (pokemon23):

Saifoo :)

OpenStudy (anonymous):

8b-4b+8=24 4b=16 b=4

OpenStudy (saifoo.khan):

8b - 4b + 8 = 24 4b = 24 - 8 4b = 16 b = 16/4 b = 4

OpenStudy (saifoo.khan):

Pokemon! :)

OpenStudy (anonymous):

8b - 4b + 8 = 24; 4b = 16; b = 4

OpenStudy (anonymous):

i like your username! pokemon

OpenStudy (pokemon23):

ty I like soccer as well soccer rules

OpenStudy (anonymous):

thanks for the medal!

OpenStudy (pokemon23):

anytime

OpenStudy (anonymous):

8b-4b+8=24 4b+8=24 ( subtract 8 from both sides ) 4b=16 (divide 4b into both sides) b=4